OpenStudy (anonymous):

2[8+4(7*2-2)] I keep getting 288, but it says the answer is 112. How would you go about getting this answer?

Mehek (mehek14):

well first you multiply 7*2

Mehek (mehek14):

then subtract the product by 2

OpenStudy (anonymous):

Did you follow the order of operations (PEMDAS)?

Mehek (mehek14):

multiply the difference by 4

Mehek (mehek14):

add 8 to the product

Mehek (mehek14):

multiply the sum by 2

Mehek (mehek14):

that's the order

OpenStudy (legends):

I got 288 also thats wierd

Mehek (mehek14):

7*2 = 14 14-2=12 12*4 = 48 48+8 = 56 56 *2 = 112

OpenStudy (anonymous):

That makes sense! Thank you so much!

Mehek (mehek14):

ur welcome ^_^